a)     A hypothetical economy is given by the following identities:

C = 3000

I = 2000

G = 2500

T = 0.2Y

MPC = 0.5

X=6500

Z=5500 + 0.2Y



               iv.           Using initial values, what will the new level of Y be if the tax rate rises to T=0.3Y?

                 v.           Calculate the budget deficit/surplus using the initial values.

               vi.           Calculate the trade balance using the initial values


1
Expert's answer
2021-07-05T08:59:40-0400

"Y=3,000+0.5(Y-0.2Y)+2,000+2,500+6,500-5,500-0.2Y\\Y=3,000+0.4Y+2,000+2,500+6,500-5,500-0.2Y\\ Y=14,000-5,500+0.2Y\\ Y=8,500+0.2Y\\ 0.8Y=8,500\\ Y=10,625"



iv)

The rise in the tax rate to 0.3Y would result in the equilibrium level of value of income to be as follows:


"Y=3,000+0.5(Y-0.3Y)+2,000+2,500+6,500-5,500-0.2Y\\

Y=3,000+0.35Y+2,000+2,500+6,500-5,500-0.2Y\\

Y=14,000-5,500+0.15Y\\

Y=8,500+0.15Y\\

0.85Y=8,500\\

 Y=10,000"



v)

The budget deficit would be the difference between the tax and the government expenditure and thus 

Budget Surplus=Tax-Government Expenditure

            "=0.3\times (10,000)-2,500\\

 =3,000-2,500\\

 =500"


vi)

The trade balance would result in the net exports

Trade Balance=Exports-Imports

           "=6,500-5,500-0.2Y\\

 =6,500-5,500-0.2\times(10,000)\\

 =6,500-5,500-2,000\\

 =6,500-7,500\\

 =-1,000"

Thus, trade balance would be a trade deficit and a budget surplus.